The perimeter of a rectangle is 36. The ratio of the lengths of the sides is 2:7. What are the lengths of the sides?
Answer by rfer(16322)

2*2x+2*7x=36
4x+14x=36
18x=36
x=36/18
x=2
2x=4
7x=14